The mass of a cyclist together with the bike is 90 kg.Calculate the increase in kinetic energy if the speed increases from 6 km/h to 12 km/h

Grade:11

1 Answers

Divyansh sharma
7 years ago
Change 6km/h to m/s =6*1000/60*60=5/3Change 12km/h to m/s=12*1000/60*60=10/3Now increase in k.E =1/2 mv²-1/2mu²=1/2m(v²-u²)=1/2*90(75/9)=1/2*750=375 Joule answer